OpenStudy (anonymous):

which function transforms the graph of y=x^2 so that it is first shifted down 2 units and then reflected across the x axis?

OpenStudy (anonymous):

a. y = -x^2 - 2 b. y = -(x+2)^2 c. y = -x^2 + 2 d. y = (-x)^2 +2

OpenStudy (anonymous):

You simply minus 2 to x^2 to shift graph down two, then make x^2 negative to flip it over the x-axis.  A would be correct.